OpenStudy (anonymous):

Suppose a man that is 6 feet tall is standing near a street light that is 10 feet high. If the shadow that is cast extends 20 feet away from the night, how far away from the light is the man?

OpenStudy (anonymous):

can you sketch the figure?

OpenStudy (anonymous):

Let d is the distance from the man to the light Apply similar triangle: ( 20 - d ) / 6 = 20 / 10 ->20 - d = 12 d = ...